Shift report 20030303 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Mar 03 2003 - 16:02:22 EST

  • Next message: Apache: "Shift report 20030304 16:00-24:00"
    ------------------------------------------------------------
    Date: Mon Mar  3 16:02:21 2003, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: 
    Log book#14 pages 44 - 45
    
    Shift Summary:
    ------------------------------------------------------------
    1. We had a beam about 4 hours.
       From 10:45 - 14:45
    
    2. MRS 40 degree, 6B
       FS   4 degree, 1/2A
       -> Run 8184 - Run 8189 : 370K trigger 3
       
       MRS 40 degree, 12B
       FS   4 degree, 1/2A
       -> Run 8190 - Run 8191 : 200K trigger 3
          Run 8190 ( D5 tripped at some point)
     
       -> 100k trigger 2 (Run 8184 - Run 8191)
    
       Pedestal, DVM Runs taken : Run 8181 - Run 8183
    
    3. Terminal Server Rebooted
       We had a problem to change polarity for D3.
       Working on T3, swiched module. (details : see E-log)
    
    Plan : MRS 40 deg 1050A - 3000k trigger 3
           FS   4 deg  1/2B -  100k trigger 2    
                   
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Mon Mar 03 2003 - 16:06:38 EST